OpenStudy (anonymous):

please help The original height of Khafre's pyramid, next to the Great Pyramid, was about 471 ft. Each side of its square base was about 708 ft. To the nearest square foot, what is the lateral area of a pyramid with those dimensions?

OpenStudy (anonymous):

Do you know the formula for lateral area?

OpenStudy (anonymous):

It is: \[1/2 \times perimeter of base \times slant height\]

OpenStudy (anonymous):

To find the perimeter of the square base, you simply add up 708ft 4 times since a square has 4 sides.

OpenStudy (anonymous):

To find the slant height, you use Pythagoras Theorem. The height is given as 471ft. The base of right triangle would be (708/2)=354

OpenStudy (anonymous):

So the slant height would be \[l^2=471^2+354^2\] l=589.2ft

OpenStudy (anonymous):

So our final answer would be: \[1/2\times 2832\times 589.2=834307.6\] Rounding this off, your answer would be 800,000 sq ft.

OpenStudy (anonymous):



OpenStudy (anonymous):

Is it clear?

OpenStudy (anonymous):

yes thank you